OpenStudy (anonymous):

how do you solve sin (5pi/3) without a calculator?

OpenStudy (chaise):

sin(5pi/3)=sin(300)=sin(60)=sqrt(3)/2 sin300=sqrt(3)/2 Make sense?

OpenStudy (anonymous):

unit circle

OpenStudy (chaise):

Oops. That should be -sin(60), and therefore -sqrt(3)/2

OpenStudy (anonymous):

why???

OpenStudy (anonymous):

do u know the unit circle

OpenStudy (anonymous):

http://www.mathpeer.com/images/trig/unit_circle.gif

OpenStudy (anonymous):

Sin 5pi/3 = Sin (pi -5pi/3) = sin -2pi/3 = - sin 2pi/3 = Sin pi -2pi/3 = sin pi/3 = sqrt(3)/2

OpenStudy (anonymous):

Use sin (pi-t) = sin t and sin-t = -sin t